OpenStudy (anonymous):

the sum of three consecutive odd numbers is at most 99. what are the numbers?

OpenStudy (anonymous):

ash, a,a+1 and a+2 are not three consecutive odd numbers.

OpenStudy (anonymous):

\( 2n+1 + 2n+3 + 2n+5 \le 99 \implies 6n \le 90 \implies 6 \le 15 \)

OpenStudy (anonymous):

so the numbers are: \( 31,33,35 \)

OpenStudy (anonymous):

foolformath, how did you get your answer?

OpenStudy (anonymous):

oh. okay. thanks.

OpenStudy (anonymous):

where'd you get the 2n?

OpenStudy (anonymous):

(2n+1) is an odd number for any \( n \in \mathbb{N} \)

OpenStudy (anonymous):

oh. okay. thank you very much!

OpenStudy (anonymous):

Glad to help :)

OpenStudy (anonymous):

foolformath, how did you get this 6n≤90 to 6≤15